# Statistical Inference Leveraging Synthetic Data with Distribution-Free Guarantees

### v1

@@ 4.3 Outlier detection with contaminated reference set
We now consider the task of conformal outlier detection (Section˜3.3) for both single and multiple testing. Conformal outlier detection guarantees error rate (Type I error/FWER) control given a reference set of pure inliers. In practice, however, one often has access to only a small inlier dataset $\mathcal{D}_{n}$—which can make conformal methods conservative (Bashari et al., 2025b)—as well as a larger, unlabeled dataset, $\tilde{\mathcal{D}}_{N}$, that is contaminated with a small percentage of outliers, say $q\%$. An ideal, but infeasible, Oracle would annotate $\tilde{\mathcal{D}}_{N}$ and use only the inliers from both datasets as reference data. As a cheap, annotation-free alternative, we use an ML model to trim the top $q\%$ of samples from $\tilde{\mathcal{D}}_{N}$ that are suspected to be outliers by the model, and treat the remainder as synthetic data consisting of “pseudo-inlier” points. Notably, this trimming can make OnlySynth less conservative, but does not guarantee error rate control at the desired level.
@@
Data. We compare the performance of conformal outlier detection methods on three benchmark tabular datasets for outlier detection: shuttle (Catlett, 1992), credit card (Group, 2013), and KDDCup99 (Stolfo et al., 1999). See Section˜B.3 for additional details on the experiments.
@@
Figure 5: Performance comparisons for outlier detection. Evaluated on three datasets: (1) shuttle, (2) credit-card, (3) KDDCup99. Left two panels: single-outlier case ($\alpha=2\%$, $\varepsilon=1\%$). Right two panels: multiple-outlier case ($\alpha=15\%$, $\varepsilon=10\%$).
Figure˜5 presents the performance for both single- and multiple-outlier testing, where both showing a similar trend. The OnlyReal method conservatively controls the error rate, but obtains low power due to the small sample size. The OnlySynth method fails to control the error rate, whereas the Oracle method achieves error rate tightly regulated around the target level, as expected. The error rate of our GESPI method is close to the nominal $\alpha$ level, while achieving substantially higher power than OnlyReal, approaching the performance of the Oracle. Additional results for both single and multiple hypothesis testing are provided in Section˜C.3.
@@ 5 Discussion
This work introduces GESPI, a general wrapper for statistical inference that safely leverages synthetic data while preserving distribution-free, finite-sample guarantees. Extensive experiments across different applications show that GESPI adapts automatically to data quality: it yields substantial gains when synthetic data are useful, but never underperforms the base method. One limitation is that the power gain depends on the quality of the synthetic data. A promising future direction is to develop adaptive methods for selecting synthetic data to further enhance statistical power.
@@ Appendix A Additional related work
Then, Theorem˜3.2 can be viewed as an upper bound on the minimax risk for the partial distribution shift problem where we observe $n$ datapoints from the original distribution, and $N$ datapoints from the shifted distribution. The algorithm $\widetilde{\mathrm{Alg}}$ certifies this. There has been work on statistical learning under a variety of distribution shifts, including the Huber contamination model (e.g., Huber, 1964, 1965, 2004; Hampel et al., 2005; Chen et al., 2016, 2018; Zhu et al., 2022, etc) and Wasserstein shifts Zhu et al. (2022); Liu and Loh (2022); Chao and Dobriban (2023). However, these works typically focus on the scenario where either (1) some random or adversarial subset of the data (that is not known to the analyst) is corrupted, or (2) all datapoints are potentially corrupted. We are not aware of studies of the scenario where a known subset is from the ground truth distribution, while the remaining subset is potentially shifted.
@@ Appendix B Experimental details
B.1 Conformal risk control for protein structure prediction
@@
Building on this, the risk from Section˜4.1
@@
For synthetic data, the true structures are unavailable. Instead, we approximate the per-residue error using the predicted alignment error (PAE) matrix. Specifically, the synthetic error for residue $i$ is taken as the mean of the $i$-th row of the PAE matrix, and we discard any synthetic protein for which more than 50% of residues exceed the 3Å threshold under this proxy.
@@
B.2 Hypothesis testing for win rate comparison between LLMs
@@
In this setting, we aim to test whether model A performs better than model B in terms of win rate. Formally, we test the null hypothesis that the win rate of model A over model B equal to $0.5$, rejecting it when we have sufficient evidence that model A wins more frequently.
In more detail, one observation corresponds to a trinomial random variable $Z\sim\text{Trinomial}(p_{\text{win}},p_{\text{equal}},p_{\text{loss}}),$ where one coordinate of $Z=(Z_{\text{win}},Z_{\text{equal}},Z_{\text{loss}})$ is equal to unity, and all others are equal to zero. After observing $n$ trials/observations, we summarize them into $N\sim\text{Trinomial}\!\left(n;\,p_{\text{win}},p_{\text{equal}},p_{\text{loss}}\right),$ where $N=(N_{\text{win}},N_{\text{equal}},N_{\text{loss}})$ is the vector of corresponding counts.
@@
Under the null hypothesis, this is a $\text{Binomial}\!\left(n-n_{\text{equal}},\tfrac{1}{2}\right)$ distribution. Hence, we can apply a randomized binomial (or sign) test conditionally on $n_{\text{equal}}$, at level $\alpha$. This test will maintain level $\alpha$ even unconditionally.
@@
Unless specified otherwise, models are run with temperature 0.6, top-p = $0.95$, top-k = $20$, and min-p = 0. For these runs, we generate 64 answers per question and take the majority vote as the final answer. The maximum token limit for all runs is 32,768.
@@
As described in the main manuscript, we consider two experimental schemes:
@@
B.3 Outlier detection with contaminated reference set
@@
For single hypothesis testing (Type I error rate control), we use the following setup: For Shuttle and KDDCup99, we use 5,000 training datapoints; for Credit Card, 2,000. Both training and reference sets are contaminated at a 5% rate. The contaminated reference set contains 2,500 datapoints, while the clean reference set, $\mathcal{D}_{n}$, contains 40 inlier points. As the outlier detection model used by the conformal outlier detection framework, we use Isolation Forest (Liu et al., 2008), implemented using scikit-learn with 100 estimators. Test sets contain 950 inliers and 50 outliers. We report the average detection power and the average Type I error over 100 independent trials.
@@ C.1 Hypothesis testing with simulated data
$\mathcal{H}_{0}:\rho=0.5\text{ versus }\mathcal{H}_{1}:\rho>0.5.$
@@
Unless stated otherwise, the real dataset contains 50 datapoints and the synthetic dataset 500 datapoints. The target Type I error level is $\alpha=5\%$, and GESPI is applied with $\varepsilon=2\%$. Under the alternative, the real data parameter is set to $\rho=0.6$, while the synthetic data parameter is $\rho_{\text{synt}}=0.55$. Each experiment is repeated 100 times to estimate the power and Type I error, and the entire procedure is repeated 100 times to evaluate variability.
@@
We begin by examining how performance varies with different choices of $\rho_{\text{synt}}$. Figure˜6 summarizes results for two regimes: $\rho=0.6$ (alternative, first row) and $\rho=0.5$ (null, bottom row).
@@
Figure 6: Performance comparison as a function of $\rho_{\text{synt}}$. Hypothesis testing methods across different values of $\rho$ applied at level $\alpha=5\%$ and $\varepsilon=2\%$. Top row: $\rho=0.6$ (alternative). Bottom row: $\rho=0.5$ (null).
Figure˜6(a) considers a setting where the alternative holds for the real data, while the synthetic data correspond to the null. Here, the synthetic data do not provide useful information for inference, and as a result, both OnlyReal and GESPI achieve comparable power.
In contrast, Figure˜6(b) presents the case where the alternative holds for both the real and synthetic data. In this setting, GESPI achieves higher power than OnlyReal, while OnlySynth attains even higher power. This is because OnlySynth naively treats the synthetic data as if it were real—a strategy that is invalid in this distribution-free setting, where the synthetic data may differ arbitrarily from the real distribution. Importantly, even when $\rho_{\text{synt}}\neq\rho$, we still observe a clear gain in power. This highlights that synthetic data do not need to perfectly match the real distribution; it suffices that the synthetic data support the same hypothesis as the real one (the alternative, in this case).
Figure˜6(c) presents the setting where the null holds for both the real and the synthetic data. All methods control the Type I error at level $\alpha$, while OnlySynth exhibits Type I error approaching zero when $\rho_{\text{synt}}<0.5$. Importantly, GESPI still benefits from the synthetic data in this setting, which highlights the point made above: the synthetic data need not perfectly match the real distribution, as long as they support the same hypothesis (here, the null).
Finally, Figure˜6(d) shows the case where the null holds for the real data, but the synthetic data follow the alternative. As in Figure˜6(a), the synthetic data are uninformative for inference. OnlySynth obtains very high Type I error and is therefore omitted from the plot (its Type I error matches the power reported in Figure˜6(b)). In contrast, GESPI controls the Type I error at most $\alpha+\varepsilon$, as guaranteed by Theorem˜F.2.
@@
Figure˜7 investigates how different values of $\varepsilon$ affects the performance of GESPI. The experiments follow a similar setup to the one in Figure˜6, with results shown as a function of $\varepsilon$ across four different scenarios: (a) the alternative holds for both real and synthetic data ($\rho=0.6$, $\rho_{\text{synt}}=0.55$), where the synthetic data provide a weaker signal against the null; (b) the null holds for both datasets ($\rho=\rho_{\text{synt}}=0.5$); (c) the alternative holds for the real data ($\rho=0.6$) and the null for the synthetic data ($\rho_{\text{synt}}=0.5$); (d) the null holds for the real data ($\rho=0.5$) and the alternative for the synthetic data ($\rho_{\text{synt}}=0.55$).
@@
Figure 7: Performance comparison as a function of $\varepsilon$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$ applied at level $\alpha=5\%$.
In scenario (a), shown in Figure˜7(a), GESPI achieves higher power than OnlyReal, and its power increases with $\varepsilon$. Intuitively, this occurs because larger values of $\varepsilon$ make the test on the real dataset at level $\alpha+\varepsilon$ more liberal, which in turn allows GESPI to rely more on the pooled-data decision: the method rejects the null if both the pooled-data test at level $\alpha$ and the real-data test at level $\alpha+\varepsilon$ reject. In scenario (b) from Figure˜7(b), where both datasets follow the null, all methods achieve Type I error close to the nominal level $\alpha$.
For scenarios (c) and (d), where the real and synthetic datasets correspond to opposing hypotheses, the synthetic data do not provide useful information for inference. In scenario (c), shown in Figure˜7(c), GESPI and OnlyReal achieve comparable power, as expected. In scenario (d), Figure˜7(d), OnlyReal controls the Type I error at level $\alpha$, while GESPI exhibits a higher Type I error, but it remains controlled at level $\alpha+\varepsilon$, as guaranteed by Theorem˜F.2.
@@
Figure˜8 compares the performance of various hypothesis testing methods as a function of the real dataset size $n$ and the synthetic dataset size $N$, under the alternative hypothesis for both real and synthetic data. Following the left panel in that figure, we can see that GESPI consistently improves power compared to the baseline, OnlyReal, across all values of $n$. The right panel demonstrates a similar trend with respect to $N$: GESPI outperforms OnlyReal, and for sufficiently large synthetic datasets ($N\geq 500$), its power remains relatively unchanged under these conditions.
Figure 8: Performance comparison as a function of the real dataset size $n$ and the synthetic dataset size $N$. Hypothesis testing methods under the alternative ($\rho=0.6$ and $\rho_{\text{synt}}=0.55$) applied at level $\alpha=5\%$ and $\varepsilon=2\%$.
@@
Figure˜9 reports performance as a function of the target Type I error $\alpha$ across the four settings considered in Figure˜7, with GESPI applied using $\varepsilon=5\%$.
@@
Figure 9: Performance comparison as a function of the target Type I error level $\alpha$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$. GESPI applied with $\varepsilon=5\%$.
In Figure˜9(a), where the alternative holds for both real and synthetic data, GESPI consistently achieves higher power than OnlyReal across all values of $\alpha$. In Figure˜9(b), where the null holds for both datasets, all methods control the Type I error close to the nominal level. As before, when the real and synthetic data correspond to opposing hypotheses (Figures˜9(c) and 9(d)), the synthetic data provide no useful information for inference. In the former case, GESPI and OnlyReal attain comparable power, while in the latter, OnlyReal controls the Type I error at level $\alpha$ and GESPI at level $\alpha+\varepsilon$, as guaranteed.
Finally, Figure˜10 presents the same analysis for $\varepsilon=2\%$, showing the same overall trends.
@@
Figure 10: Performance comparison as a function of the target Type I error level $\alpha$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$. GESPI applied with $\varepsilon=2\%$.
C.2 Conformal risk control for protein structure prediction

### v2

@@ 4.3 Multiple Hypothesis Testing for Outlier Detection
We now consider the task of conformal outlier detection (introduced in Remark˜3.1) for multiple testing. Conformal outlier detection guarantees FWER control given a reference set of pure inliers. In practice, however, one often has access to only a small inlier dataset $\mathcal{D}_{n}$—which can make conformal methods conservative (Bashari et al., 2025b)—as well as a larger, unlabeled dataset, $\tilde{\mathcal{D}}_{N}$, that is contaminated with a small percentage of outliers, say $5\%$. An ideal, but infeasible, Oracle would annotate $\tilde{\mathcal{D}}_{N}$ and use only the inliers from both datasets as reference data. As a cheap, annotation-free alternative, we use an ML model to trim the top $q\%$ of samples from $\tilde{\mathcal{D}}_{N}$ that are suspected to be outliers by the model. We then treat the remainder as synthetic data consisting of “pseudo-inlier” points. Notably, this trimming can make OnlySynth less conservative, but does not guarantee error rate control at the desired level.
@@
Data. We compare the performance of conformal outlier detection methods on three benchmark tabular datasets for outlier detection: shuttle (Catlett, 1992), credit card (Group, 2013), and KDDCup99 (Stolfo et al., 1999). See Section˜B.4 for additional details on the experiments.
Figure˜5 shows the performance of multiple testing methods. In the right panel, we report the FWER to verify the validity of GESPI. The results show trends similar to those observed across all applications: GESPI achieves FWER close to the target $\alpha$ with lower variability than Guardrail.
@@
Figure 5: Performance comparisons for outlier detection. Right: average FWER evaluated on three datasets: (1) Shuttle, (2) Credit-card, (3) KDDCup99, with $\alpha=15\%$. Left: empirical power versus empirical FWER on the Shuttle dataset. For both panels, the trimming proportion is $q=2.5\%$ and $\varepsilon=10\%$.
To provide a more detailed comparison of power, we present ROC curves (left panel) that compare the methods when they achieve the same empirical error. As portrayed, GESPI achieves higher power compared to using only real data, with performance close to that of Oracle. This shows that GESPI not only attains empirical error near the target $\alpha$ (as in the right panel) but also improves power compared to using only real data when synthetic data are of high quality. Additional results—power under FWER, single hypothesis testing, and effect of trimming proportions—are in Section˜C.6.
@@ 5 Discussion
This work introduces GESPI, a new paradigm for leveraging synthetic data in statistical inference while preserving distribution-free, finite-sample error rate control guarantees. Extensive experiments across different applications show that GESPI adapts automatically to the synthetic data quality: it yields substantial gains when synthetic data are useful, but never underperforms the base method. One limitation is that the power gain depends on the quality of the synthetic data. A promising future direction is to develop adaptive methods for selecting synthetic data to further enhance statistical power.
@@ Appendix A Additional Related Work
Then, Theorem˜3.3 can be viewed as an upper bound on the minimax risk for the partial distribution shift problem where we observe $n$ datapoints from the original distribution, and $N$ datapoints from the shifted distribution. The algorithm $\widetilde{\mathrm{Alg}}$ certifies this. There has been work on statistical learning under a variety of distribution shifts, including the Huber contamination model (e.g., Huber, 1964, 1965, 2004; Hampel et al., 2005; Chen et al., 2016, 2018; Zhu et al., 2022, etc) and Wasserstein shifts Zhu et al. (2022); Liu and Loh (2022); Chao and Dobriban (2023). However, these works typically focus on the scenario where either (1) some random or adversarial subset of the data (that is not known to the analyst) is corrupted, or (2) all datapoints are potentially corrupted. We are not aware of studies of the scenario where a known subset is from the ground truth distribution, while the remaining subset is potentially shifted.
@@
A related line of work within semi-supervised, initiated by prediction-powered inference (Angelopoulos et al., 2023a), studies how to combine a small labeled dataset with a larger set of synthetically generated labels to obtain unbiased inference and valid confidence intervals for parameters of interest. This framework has since been extended to more efficient procedures (Angelopoulos et al., 2023b), to unbiased model evaluation using synthetic labels (Boyeau et al., 2025; Fisch et al., 2024; Oosterhuis et al., 2024), and to ranking problems (Chatzi et al., 2024). Notably, as in semi-supervised settings, this line of work assumes that the covariates of the unlabeled and labeled data are i.i.d., in striking contrast to GESPI.
@@ Appendix B Experimental Details
B.1 Conformal prediction for image classification
Data.
As mentioned in the main manuscript, we adopt the experimental setup of Bashari et al. (2025a), along with the FLUX-generated images used therein. We work with a subset of 30 ImageNet classes and their corresponding synthetic images generated by FLUX (Black Forest Labs, 2024); see Bashari et al. (2025a, Appendix H.4.2) for details on the image-generation procedure. The 30 classes are listed in Tables˜S3 and S4.
Model.
Following (Bashari et al., 2025a), we use a CLIP model (Radford et al., 2021a) as the predictive model and the adaptive prediction sets (APS) score function (Romano et al., 2020); see Bashari et al. (2025a, Appendices C.1 and H.3) for additional details on the score function and predictive model.
Experimental setup.
Unless otherwise specified, for each class the real dataset contains 15 ImageNet images, and the synthetic dataset contains 1,000 images. The test set includes 500 real images per class. The target coverage level is $1-\alpha=95\%$ with $\varepsilon=2\%$. All results are averaged over 100 runs with different data splits.
B.2 Conformal Risk Control for Protein Structure Prediction
@@
Building on this, the risk from Section˜4.2
@@
For synthetic data, the true structures are unavailable. Instead, we approximate the per-residue error using the predicted alignment error (PAE) matrix. Specifically, the synthetic error for residue $i$ is taken as the mean of the $i$-th row of the PAE matrix, and we discard any synthetic protein for which more than 50% of residues exceed the 3Å threshold according to this proxy.
@@
B.3 Hypothesis Testing for Win Rate Comparison between LLMs
@@
In this setting, we aim to test whether model A performs better than model B in terms of win rate. Formally, we test the null hypothesis that the win rate of model A over model B is at most $0.5$, rejecting it when we have sufficient evidence that model A wins more frequently.
In more detail, one observation corresponds to a trinomial random variable $Z\sim\text{Trinomial}(p_{\text{win}},p_{\text{equal}},p_{\text{loss}}),$ where one coordinate of $Z=(Z_{\text{win}},Z_{\text{equal}},Z_{\text{loss}})$ is equal to unity, and all others are equal to zero. After observing $n$ independent trials/observations, we summarize them into $N\sim\text{Trinomial}\!\left(n;\,p_{\text{win}},p_{\text{equal}},p_{\text{loss}}\right),$ where $N=(N_{\text{win}},N_{\text{equal}},N_{\text{loss}})$ is the vector of corresponding counts.
@@
Under the null hypothesis, this is a $\text{Binomial}\!\left(n-n_{\text{equal}},q\right),\text{with }q\leq\tfrac{1}{2}.$ distribution. Since the binomial distribution has the monotone likelihood ratio property, it suffices to perform the test at $q=0.5$. Hence, we can apply a randomized binomial (or sign) test conditionally on $n_{\text{equal}}$, using $q=0.5$, at level $\alpha$. This test will maintain level $\alpha$ even unconditionally.
@@
For the K-way majority vote experiment, we use deepseek-ai/DeepSeek-R1-Distill-Qwen-7B with different values of $K\in\{1,4,8,16,64\}$.
Unless specified otherwise, models are run with the following settings: temperature 0.6, top-p = $0.95$, top-k = $20$, and min-p = 0. For each run, we generate $K=64$ answers per question and take the majority vote as the final answer. The maximum token limit for all runs is 32,768.
@@
For each experiment, we consider two experimental schemes:
@@
Note that for these experiments, the variability of the error depends only on the number of trials, which is the same across all methods. Therefore, unlike in the other applications presented in this paper, the variability in the error does not differ between methods.
B.4 Single and Multiple Hypothesis Testing for Outlier Detection
@@
For single hypothesis testing (Type I error rate control), we use the following setup: For Shuttle and KDDCup99, we use 5,000 training datapoints; for Credit Card, 2,000. Both training and reference sets are contaminated at a 5% rate. The contaminated reference set contains 2,500 datapoints, while the clean reference set, $\mathcal{D}_{n}$, contains 40 inlier points. We use Isolation Forest (Liu et al., 2008), implemented using scikit-learn with 100 estimators, as the outlier detection model used by the conformal outlier detection framework. Specifically, the trained model is used as a score function for computing conformal p-values. For each datapoint, it assigns a score, where a larger score indicates stronger evidence that this point may be an outlier. After computing the scores on the contaminated set, we trim the top $p\%$ of scores to attain the synthetic data. Test sets contain 950 inliers and 50 outliers. We report the average detection power and the average Type I error over 100 independent trials.
@@
In both experiments, we apply GESPI with $\varepsilon$ chosen proportional to the target level $\alpha$: we set $\varepsilon=1\%$ for single-hypothesis testing and $\varepsilon=10\%$ for multiple-hypothesis testing.
@@ C.1 Hypothesis Testing with Simulated Data
$\mathcal{H}_{0}:\rho\leq 0.5\text{ versus }\mathcal{H}_{1}:\rho>0.5.$
@@
Unless stated otherwise, the real dataset contains 50 datapoints and the synthetic dataset 500 datapoints. The target Type I error level is $\alpha=5\%$, and GESPI is applied with $\varepsilon=2\%$. Under the alternative, the real data parameter is set to $\rho=0.6$, while the synthetic data parameter is $\rho_{\text{synt}}=0.55$, illustrating a scenario where the synthetic data have a lower signal yet are still useful due to the larger sample size. Each experiment is repeated 100 times to estimate the power and Type I error, and the entire procedure is repeated 100 times to evaluate variability.
@@
We begin by examining how performance varies with different choices of $\rho_{\text{synt}}$. Figure˜S1 summarizes results for two regimes: $\rho=0.6$ (alternative, first row) and $\rho=0.5$ (null, bottom row).
@@
Figure S1: Performance comparison as a function of $\rho_{\text{synt}}$. Hypothesis testing methods across different values of $\rho$ applied at level $\alpha=5\%$ and $\varepsilon=2\%$. Top row: $\rho=0.6$ (alternative). Bottom row: $\rho=0.5$ (null).
Figure˜1(a) considers a setting where the alternative holds for the real data, while the synthetic data correspond to the null. Here, the synthetic data do not provide useful information for inference, and as a result, both OnlyReal and GESPI achieve comparable power.
In contrast, Figure˜1(b) presents the case where the alternative holds for both the real and synthetic data. In this setting, GESPI achieves higher power than OnlyReal. Its power is comparable to Guardrail, which represents the maximum power GESPI can attain, as stated in Theorems˜3.4 and F.2. Meanwhile, OnlySynth attains even higher power. This is because OnlySynth naively treats the synthetic data as if it were real—a strategy that is invalid in this distribution-free setting, where the synthetic data may differ arbitrarily from the real distribution. Importantly, even when $\rho_{\text{synt}}\neq\rho$, we still observe a clear gain in power. This highlights that synthetic data do not need to perfectly match the real distribution; it suffices that the synthetic data support the same hypothesis as the real one (the alternative, in this case).
Figure˜1(c) presents the setting where the null holds for both the real and the synthetic data. All methods control the Type I error at level $\alpha$, except Guardrail, which controls it at a higher level $\alpha+\varepsilon$. This is expected but not desired, since the goal is to control the Type I error at level $\alpha$. Note that OnlySynth exhibits Type I error approaching zero when $\rho_{\text{synt}}<0.5$. Importantly, GESPI still benefits from the synthetic data in this setting, which highlights the point made above: the synthetic data need not perfectly match the real distribution, as long as they support the same hypothesis (here, the null).
Finally, Figure˜1(d) shows the case where the null holds for the real data, but the synthetic data follow the alternative. As in Figure˜1(a), the synthetic data are uninformative for inference. OnlySynth obtains very high Type I error and is therefore omitted from the plot (its Type I error matches the power reported in Figure˜1(b)). In contrast, GESPI controls the Type I error at most $\alpha+\varepsilon$, as guaranteed by Theorem˜F.2.
@@
Figure˜S2 investigates how different values of $\varepsilon$ affects the performance of GESPI. The experiments follow a similar setup to the one in Figure˜S1, with results shown as a function of $\varepsilon$ across four different scenarios: (a) the alternative holds for both real and synthetic data ($\rho=0.6$, $\rho_{\text{synt}}=0.55$), where the synthetic data provide a weaker signal against the null; (b) the null holds for both datasets ($\rho=\rho_{\text{synt}}=0.5$); (c) the alternative holds for the real data ($\rho=0.6$) and the null for the synthetic data ($\rho_{\text{synt}}=0.5$); (d) the null holds for the real data ($\rho=0.5$) and the alternative for the synthetic data ($\rho_{\text{synt}}=0.55$).
@@
Figure S2: Performance comparison as a function of $\varepsilon$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$ applied at level $\alpha=5\%$.
In scenario (a), shown in Figure˜2(a), GESPI achieves higher power than OnlyReal, and its power increases with $\varepsilon$. Intuitively, this occurs because larger values of $\varepsilon$ make the test on the real dataset at level $\alpha+\varepsilon$ more liberal (as reflected by the higher power of Guardrail), which in turn allows GESPI to rely more on the pooled-data decision: the method rejects the null if both the pooled-data test at level $\alpha$ and the real-data test at level $\alpha+\varepsilon$ reject. In scenario (b) from Figure˜2(b), where both datasets follow the null, all methods achieve Type I error close to the nominal level $\alpha$, except for Guardrail, which by design attains the higher level $\alpha+\varepsilon$.
For scenarios (c) and (d), where the real and synthetic datasets correspond to opposing hypotheses, the synthetic data do not provide useful information for inference. In scenario (c), shown in Figure˜2(c), GESPI and OnlyReal achieve comparable power, as expected. In scenario (d), Figure˜2(d), OnlyReal controls the Type I error at level $\alpha$, while GESPI exhibits a higher Type I error, but it remains controlled at level $\alpha+\varepsilon$, as guaranteed by Theorem˜F.2.
@@
Figure˜S3 compares the performance of various hypothesis testing methods as a function of the real dataset size $n$ and the synthetic dataset size $N$, under the alternative hypothesis for both real and synthetic data. Following the left panel in that figure, we can see that GESPI consistently improves power compared to the baseline, OnlyReal, across all values of $n$. The right panel demonstrates a similar trend with respect to $N$: GESPI outperforms OnlyReal, and for sufficiently large synthetic datasets ($N\geq 500$), its power remains relatively unchanged under these conditions.
Figure S3: Performance comparison as a function of the real dataset size $n$ and the synthetic dataset size $N$. Hypothesis testing methods under the alternative ($\rho=0.6$ and $\rho_{\text{synt}}=0.55$) applied at level $\alpha=5\%$ and $\varepsilon=2\%$.
@@
Figure˜S4 reports performance as a function of the target Type I error $\alpha$ across the four settings considered in Figure˜S2, with GESPI applied using $\varepsilon=5\%$.
@@
Figure S4: Performance comparison as a function of the target Type I error level $\alpha$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$. GESPI applied with $\varepsilon=5\%$.
In Figure˜4(a), where the alternative holds for both real and synthetic data, GESPI consistently achieves higher power than OnlyReal across all values of $\alpha$, with power comparable to Guardrail. In Figure˜4(b), where the null holds for both datasets, all methods control the Type I error close to the nominal level, except for Guardrail, which controls the error at level $\alpha+\varepsilon$. Together, these figures illustrate a setting in which the synthetic data are of high quality: OnlyReal and GESPI achieve the same Type I error, while GESPI attains higher power.
We now turn to a scenario in which the synthetic data are of poor quality. As before, when the real and synthetic data correspond to opposing hypotheses (Figures˜4(c) and 4(d)), the synthetic data provide no useful information for inference. In the former case, GESPI and OnlyReal attain comparable power, while in the latter, OnlyReal controls the Type I error at level $\alpha$ and GESPI at level $\alpha+\varepsilon$, as guaranteed.
Finally, Figure˜S5 presents the same analysis for $\varepsilon=2\%$, showing the same overall trends.
@@
Figure S5: Performance comparison as a function of the target Type I error level $\alpha$. Hypothesis testing methods across different values of $\rho$ and $\rho_{\text{synt}}$. GESPI applied with $\varepsilon=2\%$.
C.2 Conformal Prediction for Image Classification
Figure˜S6 shows examples of images generated using FLUX(Black Forest Labs, 2024) for the Siberian Husky class. The images are visually high-quality and appear realistic. However, as shown in Figure˜3, these images constitute low-quality synthetic data for statistical purposes, as their distribution deviates significantly from the real data. This is evidenced by the actual coverage of OnlySynth being $\approx 60\%$, whereas the target coverage is 95%.
Figure S6: Examples of images generated using FLUX (Black Forest Labs, 2024) for the Siberian Husky class.
C.2.1 Comparing GESPI and OnlyReal
Figure˜S7 revisits the class-conditional coverage experiment presented in Figure˜3, comparing GESPI and OnlyReal when they achieve the same mean empirical coverage. Specifically, we consider multiple target levels $\alpha$ and plot performance as a function of the observed empirical coverage. Figure˜S8 shows the same setting but for the one-sided version of GESPI. Since both GESPI and OnlyReal use the same machine-learning model and the same prediction-set construction base algorithm, and obtain the same mean empirical coverage, the meaningful quantity to examine is the variance.
The variance reflects the scenario in which a practitioner would run the procedure multiple times with different realizations of the data. Accordingly, each point in the figures represents the holdout-data-conditional coverage observed from a single run using one data realization.
We observe that GESPI obtains lower variance than OnlyReal. In addition, GESPI attains coverage levels that OnlyReal cannot due to the small sample size. The achieved variance reduction demonstrates another benefit of GESPI, which goes beyond the theoretical guarantees on risk control and power. When synthetic data are of high quality, GESPI yields smaller variance and therefore higher stability. This variance reduction is important in practice, as a practitioner has access to only a single dataset, and lower variance ensures that the holdout-data-conditional coverage is more likely to be close to the target level. All in all, this experiment provides further evidence that GESPI is useful over using only the limited real data.
Figure S7: Performance of GESPI and OnlyReal as a function of the mean empirical coverage for image classification on ImageNet. Results shown for the Magpie class for varying $\alpha$ levels. GESPI applied with $\varepsilon=10\%$.
Figure S8: Performance of one-sided GESPI and OnlyReal as a function of the mean empirical coverage. Same experiment as Figure˜S7, but using the one-sided guardrail version of GESPI.
Note that the synthetic data used in this experiment are not perfect, but are of high quality. To illustrate this, Figure˜S9 presents the performance of OnlySynth and Synth+Real across varying miscoverage levels $\alpha$. These methods are intended only to illustrate the quality of the synthetic data and do not provide any distribution-free risk control guarantees. As shown, both methods consistently achieve coverage slightly below the target $1-\alpha$, indicating that the synthetic data alone does not suffice here. Nevertheless, as evidenced by Figures˜S7 and S8, the synthetic data are still of high quality, as GESPI achieves lower variance than OnlyReal.
Figure S9: Performance of OnlySynth and Synth+Real as a function of the target miscoverage rate $\alpha$ for image classification on ImageNet. Same experimental setup as Figure˜S7. The dashed black lines indicate the target coverage level $1-\alpha$. These methods illustrate the quality of the synthetic data but do not provide any distribution-free guarantees.
C.2.2 Comparison of GESPI and SPI
In this section, we compare GESPI and SPI on an image classification task using ImageNet, following the experimental setup of Figure˜3. In the conformal prediction setting, both methods are applicable. The main difference is that, unlike SPI, GESPI leverages the pooled real and synthetic data.
This distinction can be practically relevant, as generating synthetic data can be costly, and one may wish to utilize them as efficiently as possible. To illustrate this point, Figure˜S10 compares the two methods as a function of the synthetic sample size $N$. For small values of $N$, GESPI is more sample-efficient, achieving coverage closer to $1-\alpha$ and producing smaller prediction sets. As $N$ increases, the performance of the two methods becomes comparable.
Figure S10: Performance of GESPI and SPI as a function of the synthetic sample size NN for image classification on ImageNet. Results shown for the Magpie class. SPI is applied with $\beta=0.4$, and all other details follow Figure˜3.
Beyond sample efficiency, GESPI offers greater methodological flexibility than SPI. In particular, GESPI allows the use of any inference procedure that takes both $(\tilde{\mathcal{D}}_{N},\mathcal{D}_{n})$ as inputs, while the guardrail bounds still hold. This flexibility allows the practitioners to filter, modify, or adapt the synthetic data using data-dependent choices without compromising validity. In contrast, SPI is restricted to applying only the split conformal prediction algorithm solely on the synthetic data.
C.3 Conformal Risk Control for Protein Structure Prediction